OpenStudy (anonymous):

SOLVE: A. (x-3)^2=36 B. x^2-10x+25=64

OpenStudy (anonymous):

a. x - 3 = +-6 x = 9 or x = -3

OpenStudy (anonymous):

(x-3)^2=36==> x=9 or x=-3 ,

jimthompson5910 (jim_thompson5910):

(x-3)^2=36 x-3=sqrt(36) or x-3=-sqrt(36) x-3=6 or x-3=-6 x=6+3 or x=-6+3 x=9 or x=-3

OpenStudy (anonymous):

(x-5)^2=64 ===> x=13 or x=-3

OpenStudy (anonymous):

b. x -5)^2 = 64 x-5 = 8 or -8 x = 13 or -3

OpenStudy (anonymous):

the answer for a. is -3 and for b too

OpenStudy (anonymous):

yes -3 is one of the roots in both equations

OpenStudy (anonymous):

thank you jimmy :)

OpenStudy (anonymous):

yw